# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ D.1.2 Proof of Lemma 19
$\textstyle\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}},Z_{I_{k}}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{A_{i}-a}{h}\right)K\left(\frac{A_{j}-a}{h}\right)|I_{k}^{c},X_{i},X_{j},Z_{i},Z_{j}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}|X_{i},X_{j},Z_{i},Z_{j}\right]\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{A_{i}-a}{h}\right)K\left(\frac{A_{j}-a}{h}\right)$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\left(\Delta f_{i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ (41)
$\textstyle\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\left(\Delta f_{i}^{k}\right)^{2}.$
@@
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\mathbb{E}\left[\left(\Delta f_{i}^{k}\right)^{2}|I_{k}^{c}\right]$
$=\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{\|f-\hat{f}^{k}\|_{L_{2}}^{2}}{h\sigma_{Q}^{2}(a)}$
[…]
@@
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(A_{i})-\beta(a))^{2}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right||I_{k}^{c}\right]^{1/2}\cdot$
@@
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(a)}\mathbb{E}\left[(\beta(A_{i})-\beta(a))^{2}\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}|A_{i}\right]K\left(\frac{A_{i}-a}{h}\right)^{2}\right]^{1/2}\cdot$
@@
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(a)}O(h^{3/2})\|f-\hat{f}^{k}\|_{L_{2}}$
$=\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}(a)}O(n^{1/2}h),$
@@
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}}\right]$
$=\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ (42)
and hence using Hölder’s inequality with $11/p+1/q=$ 1
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$=\frac{1}{h\sigma_{Q}^{2}(a)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}|I_{k}^{c}\right]$
$\leq\frac{1}{h\sigma_{Q}^{2}(a)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]^{p}K\left(\frac{A_{i}-a}{h}\right)^{2p}\right]^{1/p}\mathbb{E}\left[|\Delta l_{i}^{k}|^{2q}|I_{k}^{c}\right]^{1/q}$
$\leq\frac{\|K\|_{\infty}^{2}}{h\sigma_{Q}^{2}(a)}\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p}}\|l-\hat{l}^{k}\|_{L_{2q}}^{2}$
@@
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(a)}\mathbb{E}\left[\left|\Delta l_{i}^{k}\Delta\phi_{i}^{k}\right|K\left(\frac{A_{i}-a}{h}\right)|I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(a)}\mathbb{E}\left[\left(\Delta l_{i}^{k}\right)^{2}|I_{k}\right]^{1/2}\mathbb{E}\left[\left(\Delta\phi_{i}^{k}\right)^{2}|I_{k}^{c}\right]^{1/2}$
$=\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(a)}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}^{k}\|_{L_{2}}$
@@ D.2 Proof or Theorem 13
By Theorem 12, $h\hat{Q}(\beta_{0}(a),a)/\sigma_{Q}(a)=O_{P}\left(h/\sqrt{Nh}\right)=o_{P}(1)$. Hence it is enough to show the following two Lemmas.
@@
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(a)(D_{i}-\phi(X_{i})))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
@@
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta_{0}(a)(D_{i}-\hat{\phi}^{k}(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$\textstyle-\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(a)(D_{i}-\phi(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\quad\cdot(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
@@ D.2.1 Proof of Lemma 20
$W_{i}=\frac{1}{\sqrt{h}\sigma_{Q}(a)}(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right).$ (43)
@@ D.2.2 Proof of Lemma 21
$\Gamma_{i}^{k}$ $=-W_{i}+\frac{1}{\sqrt{h}\sigma_{Q}(a)}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(a)(D_{i}-\hat{\phi}^{k}(X_{i})))\cdot$
$\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right).$
@@
$\Gamma_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(a)}\Bigl{[}\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{i}^{k}-\Delta f_{i}^{k}\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(a)\Delta\phi_{i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(a)\Delta\phi_{i}^{k}\right)\left(\Delta\phi_{i}^{k}-\Delta f_{i}^{k}\right)\Bigr{]}K\left(\frac{A_{i}-a}{h}\right).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (44)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (45)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (46)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (47)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta l_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (48)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (49)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (50)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (51)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\left(\Delta\phi_{i}^{k}\right)^{4}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1)$ (52)
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta\phi_{i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$ $=o_{P}(1).$ (53)
@@
$\mathbb{E}\left[\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}|I_{k}^{c},Z_{I_{k}},X_{I_{k}}\right]$
$=\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\mathbb{E}\left[\epsilon_{i}^{2}|Z_{i},X_{i}\right](\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2},$
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\left|(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right|=o_{P}(1).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\left|(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right|\right)^{2}.$
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right|=o_{P}(1).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right|\right)^{2}.$

### v2

@@ D.1.2 Proof of Lemma 18
$\textstyle\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}},Z_{I_{k}}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{V_{i}-v}{h}\right)K\left(\frac{V_{j}-v}{h}\right)|I_{k}^{c},X_{i},X_{j},Z_{i},Z_{j}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}|X_{i},X_{j},Z_{i},Z_{j}\right]\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{V_{i}-v}{h}\right)K\left(\frac{V_{j}-v}{h}\right)$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\left(\Delta f_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ (41)
$\textstyle\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\left(\Delta f_{i}^{k}\right)^{2}.$
@@
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left(\Delta f_{i}^{k}\right)^{2}|I_{k}^{c}\right]$
$=\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{\|f-\hat{f}^{k}\|_{L_{2}}^{2}}{h\sigma_{Q}^{2}(v)}$
By assertions 2 and 4 of Assumption 5, the last expression is $o_{P}(1)$. By Lemma 24 from Section D.5, we obtain (31). In exactly the same way, one can also show (32).
@@
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right||I_{k}^{c}\right]^{1/2}\cdot$
@@
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))^{2}\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}|V_{i}\right]K\left(\frac{V_{i}-v}{h}\right)^{2}\right]^{1/2}\cdot$
@@
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}O(h^{3/2})\|f-\hat{f}^{k}\|_{L_{2}}$
$=\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}(v)}O(n^{1/2}h),$
@@
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}}\right]$
$=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ (42)
and hence using Hölder’s inequality with $1/p+1/q=1$
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$=\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}|I_{k}^{c}\right]$
$\leq\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]^{p}K\left(\frac{V_{i}-v}{h}\right)^{2p}\right]^{1/p}\mathbb{E}\left[|\Delta l_{i}^{k}|^{2q}|I_{k}^{c}\right]^{1/q}$
$\leq\frac{\|K\|_{\infty}^{2}}{h\sigma_{Q}^{2}(v)}\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p}}\|l-\hat{l}^{k}\|_{L_{2q}}^{2}$
@@
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}\right|K\left(\frac{V_{i}-v}{h}\right)|I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[\left(\Delta l_{i}^{k}\right)^{2}|I_{k}\right]^{1/2}\mathbb{E}\left[\left(\Delta\phi_{2,i}^{k}\right)^{2}|I_{k}^{c}\right]^{1/2}$
$=\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(v)}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$
@@ D.2 Proof or Theorem 12
By Theorem 11, $h\hat{Q}(\beta_{0}(v),v)/\sigma_{Q}(v)=O_{P}\left(h/\sqrt{Nh}\right)=o_{P}(1)$. Hence it is enough to show the following two Lemmas.
@@
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
@@
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta_{0}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\textstyle-\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\quad\cdot(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
@@ D.2.1 Proof of Lemma 19
$W_{i}=\frac{1}{\sqrt{h}\sigma_{Q}(v)}(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right).$ (43)
@@ D.2.2 Proof of Lemma 20
$\Gamma_{i}^{k}$ $=-W_{i}+\frac{1}{\sqrt{h}\sigma_{Q}(v)}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))\cdot$
$\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right).$
@@
$\Gamma_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\Bigl[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)\Bigr]K\left(\frac{V_{i}-v}{h}\right).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (44)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (45)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (46)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (47)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta l_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (48)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (49)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (50)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (51)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (52)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1).$ (53)
@@
$\mathbb{E}\left[\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}|I_{k}^{c},Z_{I_{k}},X_{I_{k}}\right]$
$=\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\epsilon_{i}^{2}|Z_{i},X_{i}\right](\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2},$
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|=o_{P}(1).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|\right)^{2}.$
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|=o_{P}(1).$
@@
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|\right)^{2}.$